OpenStudy (anonymous):

Let f(x) = x + 9 and g(x) = x – 3. What is f(x) • g(x)?

Parth (parthkohli):

\[\rm f(x) = x + 9 , \; g(x) = x - 3\]So,\[\rm f(x) + g(x) \qquad \Rightarrow \qquad (x + 9)(x - 3)\]FOIL...

Parth (parthkohli):

Can you do that?

OpenStudy (anonymous):

@ParthKohli yes, one sec let me work it out

OpenStudy (anonymous):

i got